OpenStudy (anonymous):

What is the value of x? A. 22.5 B. 45 C. 63.6 D. 50

OpenStudy (anonymous):

here is the image

OpenStudy (anonymous):

I believe the answer is B, 45 because.... I assumed that this was a 45-45-90 triangle so I did tan(45)=x/45 x=45 tan (45) x=45 am I right?

OpenStudy (anonymous):

that would be correct

OpenStudy (anonymous):

@Spectrum great! thank you!

OpenStudy (anonymous):

one thing though.... a 45 -45 -90 triangle deals with angles..and those are 45 side lengths ..so its not bc of that..:)

OpenStudy (anonymous):

@Spectrum oh.. well I assumed that it was a 45-45-90 triangle because the angle on the top is split into two so I did 90/2 is 45 which makes 45 degrees..?

OpenStudy (mathstudent55):

Your calculations are correct, but there's no need to assume it's a 45-45-90 triangle, and you shouldn't assume that. You should prove it, which is doable.

OpenStudy (anonymous):

but how can I prove that this is a 45-45-90 triangle? I just did 90/2

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The triangles are congruent, so corresponding parts are congruent.

OpenStudy (anonymous):

yes ok

OpenStudy (mathstudent55):